TREX: Logbook Entries

TREX: TRH Messages: 4 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
252 Thu 27-Apr-2006T/RH 30m West restartedwestpoulos
193 Sat 15-Apr-20065m TRH filter checkcentralmilitzer
55 Wed 15-Mar-2006Service visit to South towersouthhorst
52 Tue 14-Mar-200615 m TRH failed at south sitesouthhorst

252: TRH, Site west, Thu 27-Apr-2006 14:41:19 PDT, T/RH 30m West restarted
The T/RH sensor at West 30 m became uni-valued at approximately noon local 26 Apr 2006. With the help of S. Oncley it was master restarted electronically at ~ 11am local 27 Apr (through rserial) to its former level of performance.
193: TRH, Site central, Sat 15-Apr-2006 12:38:43 PDT, 5m TRH filter check
15Apr06, 10:30PDT
The 5m TRH was off for a few minutes while inspecting its filter.
Good news, it appeared very clean although there was a bit of this fine dust on the support piece.
If this site's TRH is clean then almost certainly the rest are also.
55: TRH, Site south, Wed 15-Mar-2006 14:33:46 PST, Service visit to South tower
March 15, ~9.30 am - ~10:30 am

Steve and Tom visited South tower this morning.
Replaced blown fuse on NIDS serial channel 8 - 15m TRH.  Replaced fuse but Stevefound the fan was not working - probably blew the fuse again.  (Turned off 
station to replace fuse).  Replaced TRH with s/n 702(?) and replaced fuse
again - fan is working.  However, after return to trailer, getting bad data
from TRH:  both T and RH reading > 200 and s/n reporting as 704.

Steve cleaned KH2O at 30 m (~9:55 am) and 5 m (~10:30 am).
Thus on tower from ~9:45 am until 10:40 am PST.

Apparently, cleaning the KH2O's decreased the voltage at 30 m from 1.5 V to 
1.3 V!

52: TRH, Site south, Tue 14-Mar-2006 15:54:05 PST, 15 m TRH failed at south site
15m TRH failed at south site this morning at 8:20 am local time, March 14.
It had also been bad (210 deg = overrange?) for some previous period, at least
since 00Z 3/12.

Brad and Tom visited site ~3 pm and recycled power to no avail.